# Self-Labeling the Job Shop Scheduling Problem

### v2

@@ 6.2 Comparison with Proximal Policy Optimization
Section 6.1 proved the quality of the SPN when trained with our self-labeling strategy. As Proximal Policy Optimization (PPO) is extensively used to train neural algorithms for scheduling problems (see Section 2), we contrast the SPN’s performance when trained with PPO and our Self-Labeling strategy (SL-$\beta$ with $\beta\in\{32,64,128\}$). For this evaluation, we adopted the same hyper-parameters of [49] for both PPO and SL by training on 40000 random instances of shape 30x20 (same training of L2D). We only double the batch size to 8 for making the training less noisy. Figure 2 reports the validation curves obtained by testing the SPN on Taillard’s benchmark every 100 training steps. We test by sampling 128 solutions and we also include the randomized results of L2D (dashed line) as a baseline for comparisons. From Figure 2, we observe that training with our self-labeling strategy results in faster convergence and produces better final models.
We justify this by hypothesizing that the reward received from partial solutions (makespan increments) may not always provide reliable guidance on how to construct the best final solution. While constructing a solution, the critical path may change with implications on past rewards. Our self-labeling strategy does not rely on partial rewards and may avoid such a source of additional “noise". Note that this is only an intuition, proving it would require a deeper analysis, which we refer to in future works. Finally, we stress that we do not claim our self-labeling strategy is superior to PPO. Instead, we believe it offers an alternative that provides a fresh perspective and potential for integration with existing RL algorithms to advance the neural combinatorial optimization field.
Figure 3: The performance of the SPN when trained by sampling $\beta$ solutions. For each shape, we report the average PG on instances of both benchmarks by sampling $512$ solutions during testing.
Figure 4: The performance of the SPN for varying numbers of sampled solutions $\beta$ at test time. For each shape, we report the average PG on instances of both benchmarks.
@@ 6.3 The effect of $\beta$ on Training
As our self-labeling strategy is based on sampling, we assess the impact of sampling a different number of solutions $\beta$ while training. We retrain a new SPN as described in Section 5.1 with a number of solutions $\beta\in\{32,64,128,256\}$, where we stop at 256 as the memory usage with larger values becomes impractical. We test the resulting models by sampling 512 solutions on all the instances of both benchmarks for a broader assessment. Figure 3 reports the average PG (the lower the better) of the trained SPN (colored markers) on each shape. To ease comparisons, we also include the results of CL (dashed line), the second-best constructive proposal in Table 2, and the MIP (dotted line).
Overall, we see that training by sampling more solutions slightly improves the model’s overall performance, as outlined by lower PGs for increasing $\beta$. We also observe that such improvement is less marked on shapes seen in training, such as in $15\times 15,20\times 15$, and $20\times 20$ shapes, while is more marked on others. This suggests that training by sampling more solutions results in better generalization, although it is more memory-demanding. However, the trends observed for the SPN in Table 2 remain consistent with smaller $\beta$, demonstrating the robustness of our self-labeling strategy.
@@ 6.4 The effect of $\beta$ on Testing
We also assess how the number of sampled solutions $\beta$ impacts the performance of our SPN at test time. To evaluate such an impact, we plot in Figure 4 the average PG on different shapes for varying $\beta\in\{32,64,128,256,512\}$. For this analysis, we use the SPN trained by sampling 256 solutions, the same used in Table 2. As in Section 6.3, we also report the results of CL and MIP to ease the comparison.
Despite the reduced number of solutions, the SPN remains a better alternative than CL, one of the best RL proposals, and still outperforms the MIP on medium and large instances. Not surprisingly, by sampling more solutions the performance of the SPN improves, but this also increases the model execution times. Although we verified that sampling more than 512 solutions further improves results, we decided to stop at $\beta=512$ as a good trade-off between performance and time. We refer the reader to Appendix F for other timing considerations.
@@ 7 Limitations
Despite the proven effectiveness of our self-labeling strategy, it is important to note that only one of the sampled solutions per training instance is used to update the model’s parameters. This approach may be sub-optimal from an efficiency standpoint. Therefore, we see significant potential in hybridizing our self-labeling strategy with existing (RL) methods to mitigate this inefficiency. Moreover, sampling multiple solutions during training requires substantial memory, which can limit batch sizes. Although we have shown that our self-labeling strategy remains effective with a small number of sampled solutions (see Section 6.3), increasing the number of samples accelerates training convergence and improves the quality of the resulting model. Consequently, developing new sampling strategies that can construct high-quality solutions without generating numerous random ones is a promising future direction. Such advancements could reduce memory usage and further enhance our methodology.
@@ 8 Conclusions
This work outlines a novel self-labeling strategy to train generative models for the Job Shop and potentially other combinatorial problems. Our contribution is twofold: an efficient model capable of generating multiple parallel solutions, and the training strategy that solely relies on model-generated solutions. Despite its simplicity, our methodology significantly outperforms PDRs and many learning proposals, even surpassing a powerful MIP solver. However, due to its constructing nature, it still lags behind state-of-the-art methods like constraint programming solvers, which nevertheless require more time. We also proved the resiliency of our approach to different parameters and architectures.
More broadly, while the right training strategy and architecture are crucial, we emphasize the paramount importance of providing the model with the right features and state representation to create effective neural algorithms for combinatorial optimization problems. A common trend in the literature is to let the model learn its own representation by applying it to the limited information contained in the problem instance. However, our experience suggests that incorporating additional problem-specific information, like the features described in Table 1 for the JSP, can significantly enhance performance, often more than adopting variations in the training strategies.
In the future, we intend to apply our methodology to other combinatorial problems and explore combinations of our self-supervised strategy with existing RL approaches.
@@ Appendix A Disjunctive Graph Features
Following ML works reviewed in Section 2, we augment the standard disjunctive graph representation of a JSP instance by incorporating additional features for each vertex $i\in V$. Table 4 details these features. Some features, such as features 1-3, have been adapted from previous works [49, 38], while others have been introduced by us to better model JSP concepts like machines and jobs. For instance, features 4-6 take the same value for operations of the same job, emphasizing which operations belong to different jobs, while features 7-9 do the same for operations to be performed on the same machine.
Table 4: The features $x_{i}\in\mathbb{R}^{15}$ associated with a vertex $i\in V$ of the disjunctive graph $G$ provide information about operation $i$ in the instance. Recall that $\tau_{i}$ represents the processing time of operation $i$, and $\mu_{i}$ denotes the machine on which the operation is performed.
@@ Appendix B Additional Comparisons with Neural Algorithms
We additionally consider learning proposals reviewed in Section 2, which were tested only on a few benchmark instances or which could not be included in Table 2 due to how results are disclosed.
Regarding proposals not included in Table 2, we highlight that our SPN outperforms the Graph Neural Network in [38] which exhibits a significantly higher average gap of 19.50% on Taillard’s benchmark. Note that only qualitative results are disclosed in [38], we did our best to report them. Similarly, we remark that the SPN’s performance is also superior to the proposal in [44] (hCP), mixing RL and constraint programming. hCP achieves an average makespan of 2670 and 5701 on Taillard and Demirkol benchmarks respectively, while our greedy SPN yet obtains an average of 2642 and 5581.
As some other learning proposals like [20, 11] were tested only on certain benchmark instances, we provide an additional evaluation on such instances for a fair comparison. We include in Table 5 the average PGs of greedy constructive approaches on the two Taillard’s instances of each shape used in the Transformer (TRL) proposal of [11] and the Deep Q-Network (DQN) of [20].
@@
SPT MWR MOR INSA TRL L2D CL DQN SPN
17.50 18.76 15.21 15.45 35.37 22.34 15.16 7.11 15.14
29.69 24.52 26.83 15.25 32.10 26.83 17.85 13.87 13.99
27.76 22.05 23.10 16.70 28.29 30.27 16.16 17.88 17.37
39.45 22.79 23.92 21.13 36.41 31.47 20.29 16.13 17.06
30.03 27.70 27.51 19.17 34.66 38.26 22.15 21.78 21.13
30.14 23.78 26.16 12.42 31.85 24.58 15.61 17.69 11.82
24.26 18.37 18.62 22.96 28.03 28.62 14.19 19.50 14.42
14.23 8.21 8.54 15.19 17.99 12.47 5.50 9.52 4.96
26.63 20.77 21.24 17.28 30.59 26.85 15.86 15.44 14.49
On this subset of instances, we see that TRL is the worst-performing approach and that DQN roughly aligns with CL. As our SPN is always the best or second best algorithm on each shape, it achieves the lowest overall average gap (Avg).
@@ Appendix C Performance on Lawrence’s Benchmark
Lawrence’s benchmark has been extensively used in the JSP literature to develop various resolution approaches, e.g. [1, 35, 50]. This benchmark includes 40 instances of 8 different shapes and features smaller (and easier) instances compared to the Taillard’s and Demirkol’s ones. To complement the evaluation of Section 5.2 on small instances, we provide in Table 6 a comparison of approaches included in Table 2 on Lawrence’s benchmark. Note that we could not include CL and NLSA as they were not tested on these instances.
Table 6: The average PGs of coded algorithms on Lawrence’s benchmark. For each row, we highlight in blue and bold the lowest (best) constructive and non-constructive gap, respectively.
@@
3.43 1.83 0.0 0.0
Table 6 shows that our SPN remains the best constructive approach for small instances. However, we observe that L2S5k achieves slightly better results than SPN512 at the cost of significantly longer execution times. Specifically, L2S5k always requires more than 70 sec on such instance shapes (as reported in [50]), whereas our SPN512 always completes in less than 0.5 sec (see also Appendix. F for timing considerations). Finally, we highlight that on small instances like those in Lawrence’s benchmark, exact methods such as MIP and CP can effectively solve them to optimality.
@@ Appendix D Statistical Analysis
Multiple (Randomized) Constructive
MOR CL SPN128 SPN512
12.53± $\,\pm\,2.6$  9.02± $\,\pm\,2.0$  7.24± $\,\pm\,1.6$  6.52± $\,\pm\,1.3$
16.38± $\,\pm\,1.5$ 10.58± $\,\pm\,1.6$  9.31± $\,\pm\,1.8$  8.81± $\,\pm\,1.4$
14.69± $\,\pm\,1.7$ 10.87± $\,\pm\,1.6$  9.95± $\,\pm\,1.3$  9.04± $\,\pm\,1.0$
17.30± $\,\pm\,4.5$ 13.98± $\,\pm\,3.8$ 10.97± $\,\pm\,3.8$ 10.61± $\,\pm\,3.6$
20.44± $\,\pm\,2.2$ 16.09± $\,\pm\,2.0$ 13.43± $\,\pm\,1.6$ 12.66± $\,\pm\,1.4$
13.52± $\,\pm\,4.5$  9.32± $\,\pm\,3.8$  5.47± $\,\pm\,3.0$  4.93± $\,\pm\,2.7$
13.96± $\,\pm\,1.5$  9.89± $\,\pm\,2.2$  8.35± $\,\pm\,2.0$  7.58± $\,\pm\,2.0$
 7.11± $\,\pm\,2.9$  3.96± $\,\pm\,2.2$  2.32± $\,\pm\,1.1$  2.06± $\,\pm\,1.2$
14.49± $\,\pm\,4.6$ 10.46± $\,\pm\,4.2$ 8.38± $\,\pm\,3.8$ 7.78± $\,\pm\,3.6$
@@
 7.74± $\,\pm\,3.0$  6.2± $\,\pm\,1.4$  0.07± $\,\pm\,0.2$ 0.1± $\,\pm\,0.1$
12.16± $\,\pm\,1.7$  8.3± $\,\pm\,1.7$  3.22± $\,\pm\,1.9$ 0.2± $\,\pm\,0.4$
11.54± $\,\pm\,1.5$  9.0± $\,\pm\,2.2$  2.88± $\,\pm\,1.6$ 0.7± $\,\pm\,0.7$
14.13± $\,\pm\,3.5$  9.0± $\,\pm\,3.0$ 10.66± $\,\pm\,3.1$ 2.1± $\,\pm\,2.2$
16.35± $\,\pm\,2.4$ 12.6± $\,\pm\,2.1$ 13.18± $\,\pm\,2.9$ 2.8± $\,\pm\,1.5$
11.01± $\,\pm\,4.3$  4.6± $\,\pm\,3.1$ 12.25± $\,\pm\,3.9$ 0.0± $\,\pm\,0.0$
11.25± $\,\pm\,2.1$  6.5± $\,\pm\,1.9$ 13.59± $\,\pm\,4.1$ 2.8± $\,\pm\,1.9$
 5.90± $\,\pm\,2.3$  3.0± $\,\pm\,1.7$ 10.97± $\,\pm\,3.5$ 3.9± $\,\pm\,1.5$
11.26± $\,\pm\,4.1$ 7.40± $\,\pm\,3.7$ 8.36± $\,\pm\,5.9$ 1.6± $\,\pm\,1.9$
Due to space constraints, we evaluate algorithms in Table 2 solely in terms of average gaps. This approach is justified as the average trends are consistent with other statistical measures. For example, Table 7 reports the average and standard deviation of the best four multiple (randomized) constructive and non-constructive approaches on Taillard’s benchmark. We omit Demirkol’s instances since not all algorithms report results for them. Each row in Table 7 presents the average gap and standard deviation for instances of a specific shape, but the last row, which reports the average and standard deviation on all instances, regardless of their shapes. This table confirms that the observations made for Table 2 remain valid when considering standard deviations. Specifically, SPN512 remains the best constructive approach, aligning closely with the RL-enhanced metaheuristic L2S5k. The MIP still has lower performance than SPN512 while CP produces the best average gaps and standard deviations.

### v3

@@ 6.2 Comparison with Proximal Policy Optimization
The previous Sec. 6.1 proved the quality of our GM when trained with SLIM. As Proximal Policy Optimization (PPO) is extensively used to train neural algorithms for scheduling problems (see Sec. 2), we contrast the GM’s performance when trained with PPO and our self-labeling strategy (SLIM-$\beta$ with $\beta\in\{32,64,128\}$). For this evaluation, we adopt the same hyper-parameters of [54] for both PPO and SLIM by training on 40 0004000040\,00040 000 random instances of shape $30\times 20$ (same training of L2D [54]). We only double the batch size to $8$ to reduce training noise. Fig. 2 reports the validation curves obtained by testing the GM on Taillard’s benchmark every $100$ training steps. We test by sampling $128$ solutions and we also include the randomized results of L2D (dashed line) as a baseline for comparisons. From Fig. 2, we observe that training with our self-labeling improvement method results in faster convergence and produces better final models.
We justify this by hypothesizing that the reward received from partial solutions (makespan increments) may not always provide reliable guidance on how to construct the best final solution. While constructing a solution, the critical path may change with implications on past rewards. Our self-labeling strategy does not rely on partial rewards and may avoid such a source of additional “noise". Note that this is only an intuition, proving it would require a deeper analysis, which we refer to in future works. Finally, we stress that we do not claim SLIM is superior to PPO. Instead, we believe it offers an alternative that provides a fresh perspective and potential for integration with existing RL algorithms to advance the neural combinatorial optimization field.
Figure 3: The GM performance when trained by sampling varying number of solutions $\beta$. For each shape, we report the average PG on instances of both benchmarks by sampling $512$ solutions during testing.
Figure 4: The GM performance (trained as in Sec. 5.1) for varying numbers of sampled solutions $\beta$ at test time. For each shape, we report the average PG on instances of both benchmarks.
@@ 6.3 The Effect of $\beta$ on Training
As our self-labeling improvement method is based on sampling, we assess the impact of sampling a different number of solutions $\beta$ while training. We retrain a new GM as described in Sec. 5.1 with a number of solutions $\beta\in\{32,64,128,256\}$, where we stop at $256$ as the memory usage with larger values becomes impractical. We test the resulting models by sampling $512$ solutions on all the instances of both benchmarks for a broader assessment. Fig. 3 reports the average PG (the lower the better) of the trained GM (colored markers) on each shape. To ease comparisons, we also include the results of CL (dashed line) – the second-best ML proposal in Tab. 2 – and the MIP (dotted line).
Overall, we see that training by sampling more solutions slightly improves the model’s performance, as outlined by lower PGs for increasing $\beta$. We also observe that such improvement is less marked on shapes seen in training, such as in $15\times 15,20\times 15$, and $20\times 20$ shapes, and more marked on others. This suggests that training by sampling more solutions results in better generalization, although it is more memory-demanding. However, the GM’s trends observed in Tab. 2 remain consistent with smaller $\beta$, proving the robustness of SLIM to variations in the number of sampled solutions $\beta$.
@@ 6.4 The Effect of $\beta$ on Testing
We also assess how the number of sampled solutions $\beta$ impacts the GM performance at test time. To evaluate such an impact, we plot in Fig. 4 the average PG on different shapes for varying $\beta\in\{32,64,128,256,512\}$. For this analysis, we use the GM trained by sampling 256 solutions, the one of Tab. 2. As in Sec. 6.3, we also report the results of CL and MIP to ease the comparison.
Despite the reduced number of solutions, the GM remains a better alternative than CL – the best RL proposal – and still outperforms the MIP on medium and large instances. Not surprisingly, by sampling more solutions the GM performance keeps improving at the cost of increased execution times. Although we verified that sampling more than $512$ solutions further improves results (see bottom of App. F), we decided to stop at $\beta=512$ as a good trade-off between performance and time. We refer the reader to App. F for other timing considerations.
6.5 Self-Labeling the Traveling Salesman Problem
Figure 5: Validation curves obtained by training with SLIM and POMO on random TSP instances with $20$ nodes. POMO20 is the best model produced in [30], trained on instances with $20$ nodes.
Finally, we provide a brief analysis to demonstrate the broader applicability of our self-labeling improvement method to other combinatorial problems. Thus, we evaluate SLIM on the Traveling Salesman Problem (TSP), a cornerstone in neural combinatorial optimization research [5, 50, 35], by using it to train the well-known Attention Model [27] on TSP instances with 20 nodes. To assess SLIM’s performance, we compare it against the established Policy Optimization with Multiple Optima (POMO) approach [30]. We train the Attention Model with both SLIM and POMO, using the same hyper-parameters and sampling strategy outlined in [30]. Training is carried out over 100,000 steps (equivalent to 1 epoch in [30]) by generating batches of 32 instances at each step.
Fig. 5 presents the validation curves of SLIM and POMO obtained by testing the model every 100 steps on small instances (with up to 100 nodes) from the TSPLIB [41]. The results are reported in terms of the average optimality gap and, as a baseline for comparison, we include the performance of the best model in [30] (POMO20) trained for hundreds of epochs. As shown, SLIM achieves faster convergence than POMO and produces a model comparable to POMO20 after just one epoch, thereby demonstrating that SLIM can be effective in other combinatorial problems.
@@ 7 Limitations
Despite the proven effectiveness of SLIM, it is important to note that only one of the sampled solutions per training instance is used to update the model. This approach may be sub-optimal from an efficiency standpoint. Therefore, we see significant potential in hybridizing our self-labeling strategy with existing (RL) methods to mitigate this inefficiency. Moreover, sampling multiple solutions during training requires substantial memory, which can limit batch sizes. Although we have shown that SLIM remains effective with a small number of sampled solutions (see Sec. 6.3), increasing their number accelerates training convergence and improves the resulting model. Consequently, developing new strategies that can sample higher-quality solutions without generating numerous random ones is a promising future direction. Such advancements could reduce memory usage and further enhance our methodology as well as others in the literature.
@@ 8 Conclusions
The key contribution of this work is the introduction of SLIM, a novel Self-Labeling Improvement Method to train generative models for the JSP and other combinatorial problems. Additionally, an efficient encoder-decoder architecture is presented to rapidly generate parallel solutions for the JSP. Despite its simplicity, our methodology significantly outperformed many constructive and learning algorithms for the JSP, even surpassing a powerful MIP solver. However, as a constructive approach, it still lags behind state-of-the-art approaches like constraint programming solvers, which nevertheless require more time. We also proved the robustness of SLIM across various parameters and architectures, and its generality by applying it successfully to the Traveling Salesman Problem.
More broadly, self-labeling might be a valuable training strategy as it eliminates the need for optimality information or the precise formulation of a Markov Decision Process. For instance, given a designed generative model (i.e., a constructive algorithm whose decisions are taken by a neural network), this strategy can be applied as-is to combinatorial problems with unconventional objectives or a combination of objective functions. In contrast, RL approaches require the careful definition of a meaningful reward function, which is often a complex and challenging task.
In future, we intend to apply our methodology to other shop scheduling and combinatorial problems as well as explore combinations of our self-supervised strategy with other existing learning approaches.
Appendix
@@ Appendix A Disjunctive Graph Features
Following ML works reviewed in Sec. 2, we augment the standard disjunctive graph representation of a JSP instance by incorporating additional features for each vertex $i\in V$. Tab. 4 details these features. Some features, such as features 1-3, have been adapted from previous works [54, 39], while others have been introduced by us to better model JSP concepts like machines and jobs. For instance, features 4-6 take the same value for operations of the same job, emphasizing which operations belong to different jobs, while features 7-9 do the same for operations to be performed on the same machine.
Note that these features as well as those outlined in Tab. 1 can also be used for other shop scheduling problems, including Flow Shop, Flexible Flow Shop, and Flexible Job Shop.
Table 4: The features $x_{i}\in\mathbb{R}^{15}$ associated with a vertex $i\in V$ of the disjunctive graph $G$ that provide information about operation $i$ in the instance. Recall that $\tau_{i}$ represents the processing time of operation $i$, and $\mu_{i}$ denotes the machine on which the operation is performed.
@@ Appendix B Additional Comparisons with Neural Algorithms
We additionally consider learning proposals reviewed in Sec. 2, which were tested only on a few benchmark instances or which could not be included in Tab. 2 due to how results are disclosed.
Regarding proposals not included in Tab. 2, we highlight that our GM outperforms the Graph Neural Network in [39] that exhibits a significantly higher average gap of $19.5\%$ on Taillard’s benchmark. Note that only qualitative results are disclosed in [39], we did our best to report them. Similarly, we remark that the GM performance is superior to the proposal in [47] (hCP), mixing RL and constraint programming. hCP achieves an average makespan of $2670$ and $5701$ on Taillard and Demirkol benchmarks respectively, while our greedy GM yet obtains an average of $2642$ and $5581$.
As some other learning proposals like [11, 20] were tested only on certain benchmark instances, we provide an additional evaluation on such instances for a fair comparison. We include in Tab. 5 the average PGs of greedy constructive approaches on the two Taillard’s instances of each shape used in the Transformer (TRL) proposal of [11] and the Deep Q-Network (DQN) of [20].
@@
SPT MWR MOR INSA TRL L2D CL DQN GM
17.5 18.8 15.2 15.5 35.4 22.3 15.2 7.1 15.1
29.7 24.5 26.8 15.3 32.1 26.8 17.9 13.9 14.0
27.8 22.0 23.1 16.7 28.3 30.3 16.2 17.9 17.4
39.5 22.8 23.9 21.1 36.4 31.5 20.3 16.1 17.0
30.0 27.7 27.5 19.2 34.7 38.3 22.15 21.8 21.1
30.1 23.8 26.2 12.4 31.85 24.6 15.6 17.7 11.8
24.3 18.4 18.6 23.0 28.03 28.6 14.2 19.5 14.4
14.2 8.2 8.5 15.2 18.0 12.5 5.5 9.5 5.0
26.6 20.8 21.2 17.3 30.6 26.9 15.9 15.4 14.5
On this subset of instances, we see that TRL is the worst-performing approach and that DQN roughly aligns with CL. As our GM is always the best or second best algorithm on each shape, it achieves the lowest overall average gap (Avg).
@@ Appendix C Performance on Lawrence’s Benchmark
Lawrence’s benchmark has been extensively used in the JSP literature to develop various resolution approaches, e.g. [1, 36, 55]. This benchmark includes $40$ instances of $8$ different shapes and features smaller (and easier) instances compared to the Taillard’s and Demirkol’s ones. To complement the evaluation of Sec. 5.2 on small instances, we provide in Tab. 6 a comparison of approaches included in Tab. 2 on Lawrence’s benchmark. Note that we could not include CL and NLSA as they were not tested on these instances.
Table 6: The average PGs of the algorithms on Lawrence’s benchmark. For each row, we highlight in blue and bold the lowest (best) constructive and non-constructive gap, respectively.
@@
3.4 1.8 0.0 0.0
Tab. 6 shows that our GM remains the best constructive approach for small instances. However, we observe that L2S5k achieves slightly better results than GM512 at the cost of significantly longer execution times. Specifically, L2S5k always requires more than 70 sec on such instance shapes (as reported in [55]), whereas our GM512 always completes in less than $0.5$ sec (see also App. F for timing considerations). Finally, we highlight that on small instances like those in Lawrence’s benchmark, exact methods such as MIP and CP can effectively solve them to optimality.
@@ Appendix D Statistical Analysis
Multiple (Randomized) Constructives
MOR CL GM128 GM512
12.5± $\,\pm\,2.6$  9.0± $\,\pm\,2.0$  7.2± $\,\pm\,1.6$  6.5± $\,\pm\,1.3$
16.4± $\,\pm\,1.5$ 10.6± $\,\pm\,1.6$  9.3± $\,\pm\,1.8$  8.8± $\,\pm\,1.4$
14.7± $\,\pm\,1.7$ 10.9± $\,\pm\,1.6$  10.0± $\,\pm\,1.3$  9.0± $\,\pm\,1.0$
17.3± $\,\pm\,4.5$ 14.0± $\,\pm\,3.8$ 11.0± $\,\pm\,3.8$ 10.6± $\,\pm\,3.6$
20.4± $\,\pm\,2.2$ 16.1± $\,\pm\,2.0$ 13.4± $\,\pm\,1.6$ 12.7± $\,\pm\,1.4$
13.5± $\,\pm\,4.5$  9.3± $\,\pm\,3.8$  5.5± $\,\pm\,3.0$  4.9± $\,\pm\,2.7$
14.0± $\,\pm\,1.5$  9.9± $\,\pm\,2.2$  8.4± $\,\pm\,2.0$  7.6± $\,\pm\,2.0$
 7.1± $\,\pm\,2.9$  4.0± $\,\pm\,2.2$  2.3± $\,\pm\,1.1$  2.1± $\,\pm\,1.2$
14.5± $\,\pm\,4.6$ 10.5± $\,\pm\,4.2$ 8.4± $\,\pm\,3.8$ 7.8± $\,\pm\,3.6$
@@
 7.7± $\,\pm\,3.0$  6.2± $\,\pm\,1.4$  0.1± $\,\pm\,0.2$ 0.1± $\,\pm\,0.1$
12.2± $\,\pm\,1.7$  8.3± $\,\pm\,1.7$  3.2± $\,\pm\,1.9$ 0.2± $\,\pm\,0.4$
11.5± $\,\pm\,1.5$  9.0± $\,\pm\,2.2$  2.9± $\,\pm\,1.6$ 0.7± $\,\pm\,0.7$
14.1± $\,\pm\,3.5$  9.0± $\,\pm\,3.0$ 10.7± $\,\pm\,3.1$ 2.1± $\,\pm\,2.2$
16.4± $\,\pm\,2.4$ 12.6± $\,\pm\,2.1$ 13.2± $\,\pm\,2.9$ 2.8± $\,\pm\,1.5$
11.0± $\,\pm\,4.3$  4.6± $\,\pm\,3.1$ 12.3± $\,\pm\,3.9$ 3.0± $\,\pm\,0.0$
11.3± $\,\pm\,2.1$  6.5± $\,\pm\,1.9$ 13.6± $\,\pm\,4.1$ 2.8± $\,\pm\,1.9$
 5.9± $\,\pm\,2.3$  3.0± $\,\pm\,1.7$ 11.0± $\,\pm\,3.5$ 3.9± $\,\pm\,1.5$
11.3± $\,\pm\,4.1$ 7.4± $\,\pm\,3.7$ 8.4± $\,\pm\,5.9$ 2.0± $\,\pm\,1.9$
Due to space constraints, we evaluate algorithms in Tab. 2 solely in terms of average gaps. This approach is justified as the average trends are consistent with other statistical measures. For example, Tab. 7 reports the average and standard deviation of the best four multiple (randomized) constructive and non-constructive approaches on Taillard’s benchmark. We omit Demirkol’s instances since not all algorithms report results for them. Each row in Tab. 7 presents the average gap and standard deviation for instances of a specific shape, but the last row, which reports the average and standard deviation on all instances, regardless of their shapes. This table confirms that the observations made for Tab. 2 remain valid when considering standard deviations. Specifically, GM512 remains the best constructive approach, aligning closely with the RL-enhanced metaheuristic L2S5k. The MIP still has worse overall performance than GM512 while CP produces the best average gaps and standard deviations.